OpenStudy (anonymous):

write a fraction equal to one half that has a denominator of 10 then subtract the fraction from nine tenths and reduce

OpenStudy (radar):

The denominator has to be a 10 and it must also = 1/2 what multiplied by 2 = 10? If you said 5, you're right so the fraction is 5/10 so subtrtact that from 9/10

OpenStudy (anonymous):

(9/10) - (5/10) = (4/10)

OpenStudy (radar):

9-5 = 4, now we have 4/10. It says reduce. 2/5 is your final answer.

OpenStudy (anonymous):

thanks